1. Convert 172 to the following equivalences:

octal

hexadecimal

Show ALL working

1
Expert's answer
2022-03-16T15:47:47-0400

To convert decimal number 172 to octal, we follow these steps:

  1. We divide 172 by 8 keeping notice of the quotient and the remainder.
  2. We continue dividing the quotient by 8 until we get a quotient of zero.
  3. Then we just write out the remainders in the reverse order to get octal equivalent of decimal number 172.

Using the above steps, here is the work involved in the solution for converting 172 to octal number:

172 / 8 = 21 with remainder 4

21 / 8 = 2 with remainder 5

2 / 8 = 0 with remainder 2

The decimal number 172 converted to octal is therefore equal to : 254.


Here is the work involved in the solution for converting 172 to hexadecimal number:

172 / 16 = 10 with remainder 12 (C)

10 / 16 = 0 with remainder 10 (A)

The decimal number 172 converted to hexadecimal is equal to : AC.